OpenStudy (anonymous):

Consumers in a certain state can choose between three long-distance telephone services: GTT, NCJ, and Dash. Aggressive marketing by all three companies results in continual shift of customers among the three services. Each year, GTT loses 25% of its customers to NCJ and 25% to Dash, NCJ loses 20% of its customers to GTT and 20% to Dash, and Dash loses 25% of its customers to GTT and 30% to NCJ. Assuming that these percentages remain valid over a long period of time, what is each company's expected market share in the long run?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

GTT's expected market share is ____% NCJ's expected market share is _____% Dash expected market share is _____ %

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

reminds me of a markov chain

OpenStudy (amistre64):

Lets unpack the novel .... Consumers in a certain state can choose between three long-distance telephone services: GTT, NCJ, and Dash. Aggressive marketing by all three companies results in continual shift of customers among the three services. Each year, GTT loses 25% of its customers to NCJ and 25% to Dash, NCJ loses 20% of its customers to GTT and 20% to Dash, Dash loses 25% of its customers to GTT and 30% to NCJ. Assuming that these percentages remain valid over a long period of time, what is each company's expected market share in the long run? i assume its asking for the stable state?

OpenStudy (anonymous):

yes i believe it is a makov chain

OpenStudy (amistre64):

i believe the movement then looks like this each year? Rows define the movements, and Columns define who gets what G N D G .50 .25 .25 N .20 .60 .20 D .25 .30 .45

OpenStudy (amistre64):

or do i have that backwards?

OpenStudy (amistre64):

lets assume thats backwards, and that the columns have to add up to 1. which way would you set it up?

OpenStudy (amistre64):

\[\begin{pmatrix} .50&.20&.25\\ .25&.60&.30\\ .25&.20&.45 \end{pmatrix}\Huge X=\Huge X\] then X is the stable setup

OpenStudy (amistre64):

this should remind you of eugene vectors and values

OpenStudy (anonymous):

what is that correct order? are u sure its not backwards? cuz i dont think it is backwards or maybe im mistaken? im not quite sure

OpenStudy (amistre64):

ive read too many things to be certain at this point. If a had a textbook in front of me id be able to be certain. Some sites i see say columns are 1, others say rows are 1 Do you have a textbook handy?

OpenStudy (anonymous):

it is an online assignment

OpenStudy (anonymous):

that is the question :p

OpenStudy (amistre64):

even online assignments have material for resourses dont they? i got a google book that has the columns as 1s so ill feel safe with that

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

they have a recurssion set up with vectors x0, x1, x2, x3 .... unti xn is the stable state. but im not sure how they come up with the initial state vector x0

OpenStudy (amistre64):

the initial state vector seems to be composed of the probability of any one person moving to either G N or D

OpenStudy (amistre64):

arent all destinations equally likely for any random person to go to?

OpenStudy (amistre64):

http://aix1.uottawa.ca/~jkhoury/markov.htm this seems to be similar to your problem, with any luck we can glean something useful

OpenStudy (amistre64):

woohoo, i got it :)

OpenStudy (anonymous):

hmm i still dont understand how to find the market share percentage for GTT, Dash, and NCJ

OpenStudy (amistre64):

we are multiplying the original matrix to itself, over and over again GND^1 * GND^1 = GND^2 GND^1 * GND^2 = GND^3 GND^1 * GND^3 = GND^4

OpenStudy (amistre64):

after a few goes at it, you should be able to see the entries evening out into a stable set

OpenStudy (amistre64):

i plugged the matrix setup into my ti83, and i am running a few iterations at the moment

OpenStudy (anonymous):

so what would be the result from multiplying it over and over again?

OpenStudy (amistre64):

im getting a limit of say: 30.62%, 40.67%, and 28.71%

OpenStudy (amistre64):

\[\begin{pmatrix} .50&.20&.25\\ .25&.60&.30\\ .25&.20&.45 \end{pmatrix}\cdot \begin{pmatrix} .50&.20&.25\\ .25&.60&.30\\ .25&.20&.45 \end{pmatrix}=T^1\] \[T^1\cdot \begin{pmatrix} .50&.20&.25\\ .25&.60&.30\\ .25&.20&.45 \end{pmatrix}=T^2\] \[T^2\cdot \begin{pmatrix} .50&.20&.25\\ .25&.60&.30\\ .25&.20&.45 \end{pmatrix}=T^3\] ... eventually the entries even out such that \[T^n\approx T^{n+1}\]

OpenStudy (amistre64):

and of course the market share is in the form of GND since that is how i setup my matrix

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

im sure there is a faster way to work it, but im not sure what it would be. but does the process make sense now?

OpenStudy (anonymous):

yes it makes a little more sense for me now haha im just not sure how to get the percentage of each market share you know?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i can spend a few minutes and type up an html to run the matrixes

OpenStudy (anonymous):

i just need help finding/showing me how to get the percentages

OpenStudy (amistre64):

do you know how to multiply matrixes?

OpenStudy (anonymous):

i forgot how to do it thats why i need someone to show me then i will be able to remember

OpenStudy (amistre64):

lets take a simple 2x2 setup \[\begin{pmatrix} a&b\\ c&d\\ \end{pmatrix}\cdot \begin{pmatrix} 1&5\\ 3&2\\ \end{pmatrix}=\begin{pmatrix} 1a+3b&5a+2b\\ 1c+3d&5c+2d\\ \end{pmatrix}\]

OpenStudy (amistre64):

the new matrix is constructed by the inner product of Ar1 and Bc2, meaning A row 1 and B col 2 given A = R1 , R2 and B = C1 C2 the new matrix developed from A times B is: R1.C1 R1.C2 R2.C2 R2.C2

OpenStudy (amistre64):

in the case of the Markov here; A and B are the same matrix ...

OpenStudy (amistre64):

\[\begin{pmatrix} 1&5\\ 3&2\\ \end{pmatrix}\cdot \begin{pmatrix} 1&5\\ 3&2\\ \end{pmatrix}=\begin{pmatrix} 1(1)+3(5)&5(1)+2(5)\\ 1(3)+3(2)&5(3)+2(2)\\ \end{pmatrix}\]

OpenStudy (amistre64):

\[\begin{pmatrix} 1(1)+3(5)&5(1)+2(5)\\ 1(3)+3(2)&5(3)+2(2)\\ \end{pmatrix} \cdot \begin{pmatrix} 1&5\\ 3&2\\ \end{pmatrix}=\\ \begin{pmatrix} 1[(1(1)+3(5)]+3(5(1)+2(5)]&5[1(1)+3(5)]+2[5(1)+2(5)]\\ 1[1(3)+3(2)]+3[5(3)+2(2)]&5[1(3)+3(2)]+2[5(3)+2(2)]\\ \end{pmatrix}\] and keeps on going and going and going

OpenStudy (amistre64):

i just wrote this up for a 3x3

OpenStudy (amistre64):

notice that in the first question you asked, when i input: .50 .20 .25 .25 .60 .30 .25 .20 .45 and run it for 100 times we get: M1 0.3625 0.27 0.2975 0.35000000000000003 0.47 0.3775 0.2875 0.26 0.325 M2 0.323125 0.294 0.3055 0.38687499999999997 0.4275 0.398375 0.29 0.27849999999999997 0.29612499999999997 M3 0.31143750000000003 0.302125 0.30645625 0.39990624999999996 0.41355000000000003 0.4042375 0.28865624999999995 0.28432499999999994 0.28930625 M4 0.3078640625 0.30485375000000003 0.3064021875 0.4044 0.40895875000000004 0.40594843750000004 0.2877359374999999 0.28618749999999993 0.287649375 M5 0.306746015625 0.3057655 0.306303125 0.405926796875 0.4074449375 0.406464421875 0.28732718749999997 0.28678956249999993 0.28723245312499995 M6 0.30639016406249997 0.30606912812500003 0.30625256015625 0.40644073828125 0.40694520624999997 0.4066241703125 0.2871690976562499 0.28698566562499994 0.28712326953124995 M7 0.3062755041015625 0.30617002171875 0.30623193152343753 0.40661271328125 0.40678010546875 0.4066746230859375 0.28711178261718745 0.28704987281249994 0.287093445390625 M8 0.3062382403613282 0.30620350015625 0.30622425172656254 0.40667003877929686 0.40672553055468746 0.40669079034960937 0.28709172085937495 0.28707096928906245 0.2870849579238281 M9 0.3062260581513672 0.30621459851132815 0.3062215234141602 0.40668909961572264 0.40670748415859376 0.4066960245185547 0.28708484223291014 0.2870779173300781 0.2870824520672851 M10 0.3062220595570557 0.3062182754199024 0.30622057962761234 0.40669542697714844 0.4067015153220117 0.4066977311848584 0.28708251346579583 0.2870802092580859 0.2870816891875293 M11 0.3062207435404065 0.306219493088875 0.3062202583476602 0.4066975251152918 0.40669954082560844 0.4066982903740769 0.2870817313443017 0.2870809660855166 0.2870814512782629 M12 0.3062203096293371 0.30621989623093837 0.30622015006821124 0.4066982203575672 0.4066988875932388 0.40669847419484007 0.2870814700130957 0.2870812161758229 0.2870813757369487 M13 0.30622016638945593 0.3062200296780727 0.30622011380731085 0.4066984506258033 0.40669867146642474 0.4066985347550415 0.28708138298474073 0.2870812988555026 0.28708135143764774 M14 0.3062201190660738 0.306220073846197 0.30622010171407565 0.40669852686826824 0.4066985999560238 0.4066985547361469 0.28708135406565793 0.28708132619777926 0.28708134354977743 M15 0.3062201034231051 0.3062200884637481 0.3062200976917116 0.40669855210717676 0.40669857629449735 0.4066985613351403 0.2870813444697181 0.2870813352417546 0.2870813409731481 M16 0.30622009825041746 0.3062200933012122 0.30622009635617087 0.4066985604609978 0.40669856846516184 0.40669856351595657 0.28708134128858476 0.287081338233626 0.28708134012787256 M17 0.3062200965395545 0.306220094902045 0.30622009591324495 0.40669856322577846 0.40669856587448794 0.4066985642369785 0.287081340234667 0.28708133922346707 0.28708133984977663 M18 0.3062200959735998 0.30622009543178685 0.30622009576646236 0.40669856414075584 0.40669856501724416 0.4066985644754313 0.28708133988564444 0.28708133955096893 0.28708133975810635 M19 0.3062200957863622 0.3062200956070845 0.30622009571784403 0.4066985644435468 0.4066985647335839 0.4066985645543063 0.28708133977009104 0.2870813396593315 0.28708133972784966 M20 0.3062200957244132 0.306220095665092 0.30622009570174574 0.406698564543746 0.406698564639721 0.4066985645803997 0.2870813397318408 0.28708133969518707 0.28708133971785454 M21 0.30622009570391606 0.306220095684287 0.3062200956964165 0.4066985645769032 0.4066985646086617 0.40669856458903264 0.2870813397191808 0.2870813397070513 0.28708133971455085 M22 0.30622009569713393 0.3062200956906387 0.3062200956946526 0.4066985645878752 0.40669856459838416 0.406698564591889 0.28708133971499095 0.2870813397109771 0.2870813397134585 M23 0.3062200956948898 0.3062200956927405 0.3062200956940687 0.4066985645915059 0.4066985645949833 0.40669856459283404 0.2870813397136044 0.2870813397122762 0.2870813397130972 M24 0.3062200956941472 0.306220095693436 0.30622009569387554 0.4066985645927073 0.406698564593858 0.4066985645931468 0.28708133971314553 0.28708133971270605 0.28708133971297767 M25 0.3062200956939015 0.30622009569366615 0.3062200956938116 0.4066985645931048 0.4066985645934856 0.40669856459325027 0.2870813397129937 0.28708133971284827 0.28708133971293814 M26 0.3062200956938202 0.3062200956937423 0.30622009569379044 0.4066985645932364 0.40669856459336234 0.40669856459328446 0.2870813397129435 0.28708133971289534 0.2870813397129251 M27 0.30622009569379327 0.3062200956937675 0.30622009569378345 0.4066985645932799 0.40669856459332154 0.4066985645932958 0.2870813397129268 0.2870813397129109 0.28708133971292077 M28 0.3062200956937844 0.30622009569377584 0.30622009569378117 0.40669856459329423 0.40669856459330805 0.40669856459329956 0.2870813397129213 0.28708133971291605 0.2870813397129193 M29 0.30622009569378145 0.3062200956937786 0.3062200956937804 0.406698564593299 0.4066985645933036 0.4066985645933008 0.2870813397129195 0.28708133971291777 0.2870813397129188 M30 0.3062200956937805 0.30622009569377956 0.3062200956937801 0.4066985645933006 0.4066985645933021 0.40669856459330117 0.2870813397129189 0.2870813397129183 0.2870813397129187 M31 0.30622009569378017 0.30622009569377984 0.30622009569378006 0.40669856459330117 0.40669856459330167 0.4066985645933013 0.2870813397129187 0.2870813397129185 0.2870813397129186 M32 0.30622009569378006 0.30622009569377995 0.30622009569378 0.4066985645933013 0.4066985645933015 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M33 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330145 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M34 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.4066985645933014 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M35 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M36 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M37 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M38 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M39 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M40 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M41 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M42 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M43 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M44 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M45 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M46 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M47 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M48 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M49 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M50 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M51 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M52 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M53 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M54 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M55 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M56 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M57 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M58 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M59 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M60 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M61 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M62 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M63 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M64 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M65 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M66 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M67 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M68 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M69 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M70 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M71 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M72 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M73 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M74 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M75 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M76 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M77 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M78 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M79 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M80 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M81 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M82 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M83 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M84 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M85 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M86 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M87 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M88 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M89 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M90 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M91 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M92 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M93 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M94 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M95 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M96 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M97 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M98 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186 M99 0.30622009569378 0.30622009569378 0.30622009569378 0.40669856459330134 0.40669856459330134 0.40669856459330134 0.2870813397129186 0.2870813397129186 0.2870813397129186

OpenStudy (amistre64):

by the time we get down to 99 iterations; the entries of each row is the same the limit that they are approaching tells us the long run effect of the movements within the market share. 30.62% 40.67% 28.71%

OpenStudy (amistre64):

now i did 100 to show you that these value dont seem to change that much from about the 10th iteration

OpenStudy (amistre64):

i spose i could modify it for a 4x4 matrix .... for the second question.

OpenStudy (anonymous):

how do i do the second question?

OpenStudy (amistre64):

when PS^(k) = S^(k+1) this is the same concept and procedure, and the one i first thought this post expected. you want to find the condition such that:\[S^{(k+1)}\approx S^{(k)}\]

OpenStudy (anonymous):

Okay so how do i find that condition